# Client Selection for Federated Policy Optimization with Environment Heterogeneity

### v6

@@ 2.1 Federated Reinforcement Learning
We model the local learning problem of each client as a finite-state infinite-horizon discounted Markov Decision Process (MDP). Accordingly, the FRL system consists of $N$ finite MDPs $\{(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)|n\in\{1,...,N\}\}$, where $\mathcal{S}$ denotes a finite set of states, $\mu$ represents the initial state distribution, $\mathcal{A}$ is a finite set of actions, and $\gamma\in(0,1)$ is the discount factor. The transition function $P_{n}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ represents the probability that the n-th MDP transits from state $s$ to $s^{\prime}$ after taking action $a$ (Sutton and Barto, 2018). The reward function $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ gives the expected reward for taking action $a$ in state $s$, and we assume rewards are bounded and non-negative. As a result, the $n$-th MDP $\mathcal{M}_{n}$ can be represented by a 6-tuple $(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)$ sharing the same state space, action space, initial state distribution, and reward function with other clients, but with possibly different transition probabilities. A client’s behavior is controlled by a stochastic policy $\pi_{n}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which outputs the probability of taking an action $a$ in a given state $s$. Throughout this work, we consider parameterized policies $\pi^{\theta}$ where $\pi^{\theta}(a|s)$ is a differentiable function of the parameter vector $\theta$. For parameterized policies with t-indexed notation, we omit the parameter vector for notation simplicity and write5 $\pi^{t}$ and $\pi^{t}_{n}$ for the global policy $\pi^{\theta^{t}}$ and the $n$-th local policy $\pi^{\theta^{t}_{n}}$, respectively. Furthermore, we define the state-value function
@@ 3.4 Impact of Aggregation Error
where $m$ is the width of the network, $\theta=(\theta_{i},\dots\theta_{m})^{T}\in\mathcal{R}^{m(d_{s}+d_{a})}$, $w=(w_{i},\dots w_{m})^{T}\in\mathcal{R}^{md_{s}}$ denotes the input weights, and $b_{i}\in\{-1,1\},\forall i\in[m]$ is the output weight. Without loss of generality, we assume that $a\in\mathbb{R}^{d_{a}},s\in\mathbb{R}^{d_{s}}$ and denote $d=d_{s}+d_{a}$. Given arbitrary $\hat{R}_{\vartheta}<\infty$, we consider the following parameter initialization, where $\vartheta$ is a placeholder for $\theta$ and $w$:
@@
$\bar{\varepsilon}_{w}$ $=\mathcal{O}\left(R_{w}^{6/5}m^{-1/10}\hat{R}_{w}^{2/5}\right),$ (18)
$\tilde{\varepsilon}_{\theta}$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right),$ (19)
$\bar{\varepsilon}_{\theta}$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right).$ (20)
@@ 4.2 Implementation
To compute $\Delta_{n}$, we adopt the tabular maximum likelihood model (Moerland et al., 2020; Strehl et al., 2009; Ornik and Topcu, 2021; Strehl and Littman, 2008) in the implementation of FedPOHCS. Given a set of trajectories, let $C_{n}(s,a)$ denote the number of times action $a$ was taken under state $s$, $C_{n}(s,a,s^{\prime})$ represent the number of times the MDP transited from state $s$ to $s\prime$ after taking action $a$, and $r_{n}$ denote the corresponding sequence of reward received. Generally speaking, the more trajectories we have, the more accurate the modeling would be. For Mountain Cars and Hoppers, the number of trajectories is 200. For HongKongOSMs, the number of trajectories is 1000. We can estimate the transition probability $P_{n}$ of the $n$-th MDP and the reward function $\mathcal{R}$ by
@@ 5 Experiments
In this section, we introduce two federated environments for empirical evaluation and evaluate the proposed client selection method from different perspectives. All experimental results are reproducible and can be accessed from the GitHub repository: https://github.com/ShiehShieh/FedPOHCS.
@@ 5.1 Environments
HongKongOSMs consists of 10 equally weighted traffic networks, each based on one of the OSM data as shown in Figure 1. Each traffic network contains one RL-controlled and 10 IDM-controlled (Intelligent Driver Model) vehicles. The 18-dimensional observation includes headway, speed, and positional information of visible neighborhoods of the RL-controlled vehicle, and only observations for vehicles running on the same and adjacent lanes are visible to the RL-controlled vehicle. The 2-dimensional action includes acceleration and lane-changing decisions. We adopted the Eclipse Simulation of Urban MObility (SUMO) simulator to conduct this experiment.
Figure 1: OSM data of ten areas in Hong Kong. (a) Causeway; (b) Central; (c) Chai Wan; (d) Clear Water Bay; (e) Kennedy Town; (f) Kai Tak; (g) North Point; (h) Po Lam; (i) Sham Shui Po; (j) Tseung Kwan O.
@@ Appendix G Proof of Lemma 12
$u_{w^{t}}^{0}(s)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[w(0)]_{i}^{T}(s)>0\right\}[w^{t}]_{i}^{T}(s),$
$f_{\theta^{t}}^{0}(s,a)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)>0\right\}[\theta^{t}]_{i}^{T}(s,a).$
@@
For any $w\in\mathcal{B}_{R_{w}}^{0},\theta\in\mathcal{B}_{R_{\theta}}^{0},s\in\mathcal{S}$, and $a\in\mathcal{A}$, we have
@@
Proof  Given any pair of model parrameters $\theta\in\mathcal{B}_{R_{\theta}}^{0}$ and $\theta^{\prime}\in\mathcal{B}_{R_{\theta}}^{0}$,
$\mathbf{1}\left\{[\theta]_{i}^{T}(s,a)>0\right\}\neq\mathbf{1}\left\{[\theta^{\prime}]_{i}^{T}(s,a)>0\right\}$
@@
$\left|[\theta^{\prime}]_{i}^{T}(s,a)\right|\leq\left|[\theta]_{i}^{T}(s,a)-[\theta^{\prime}]_{i}^{T}(s,a)\right|\leq\left\|[\theta]_{i}-[\theta^{\prime}]_{i}\right\|_{2}.$
@@
$\left|f_{\theta}(s,a)-f_{\theta}^{0}(s,a)\right|$
$=$ $\frac{1}{\sqrt{m}}\left|\sum_{i=1}^{m}b_{i}\cdot\left(\mathbf{1}\left\{[\theta]^{T}_{i}(s,a)>0\right\}-\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)>0\right\}\right)\cdot[\theta]_{i}^{T}(s,a)\right|$
$\leq$ $\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\left|\mathbf{1}\left\{[\theta]^{T}_{i}(s,a)>0\right\}-\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)>0\right\}\right|\cdot\left|[\theta]_{i}^{T}(s,a)\right|$
$\leq$ $\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}\cdot\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}.$ (60)
@@
  • Case 1: $G_{1}=\frac{1}{\sqrt{m}}\sum_{i\in C_{1}}\left(\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}\right)\cdot\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}$, where $C_{1}=\left\{i\in[m]:\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\leq\Delta\right\}$ and $\Delta>0$.
@@
    $G_{1}$ $\leq\frac{1}{\sqrt{m}}\sum_{i\in C_{1}}\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}=\mathcal{O}\left(m^{1/2}\Delta^{2}R_{0}^{-1}\right).$ (61)
@@
  • Case 2: $G_{2}=\frac{1}{\sqrt{m}}\sum_{i\in C_{2}}\left(\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}\right)\cdot\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}$, where $C_{2}=\left\{i\in[m]:\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}>\Delta\right\}$ and $\Delta>0$.
    We have $\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}/\left\|[\theta(0)]_{i}\right\|_{2}>\Delta/\hat{R}_{\theta}$. Consequently, there exits a constant $c\geq\hat{R}_{\theta}/\Delta$ such that for any layer $i\in C_{2},a\in\mathcal{A},s\in\mathcal{S}$, it holds that
    $\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}\leq 1\leq c\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}/\left\|[\theta(0)]_{i}\right\|_{2}.$ (62)
    Next, by the Cauchy-Schwarz inequality and $\left\|\vartheta-\theta(0)\right\|_{2}\leq R_{\theta}$, we have
    $G_{2}$ $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{\sum_{i\in C_{2}}\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}}.$ (63)
@@
    $\mathbb{E}_{\text{init}}\left[G_{2}\right]$ $\leq\frac{R_{\theta}}{\sqrt{m}}\mathbb{E}_{\text{init}}\left[\sqrt{\sum_{i\in C_{2}}\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}}\right]$
    $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\mathbf{1}\left\{[\theta(0)]_{i}^{T}(s,a)\leq\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}\right\}\right]}$
    $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{c\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}/\|[\theta(0)]_{i}\|_{2}\right]}.$
@@
    $\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}/\|[\theta(0)]_{i}\|_{2}\right]$
    $\leq$ $\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}^{2}\right]^{1/2}\cdot\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\|[\theta(0)]_{i}\|_{2}^{-2}\right]^{1/2}$
    $\leq$ $R_{\theta}\cdot\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\|[\theta(0)]_{i}\|_{2}^{-2}\right]^{1/2},$ (64)
    where the second inequality follows from $\sum_{i=1}^{m}\left\|[\theta]_{i}-[\theta(0)]_{i}\right\|_{2}^{2}=\left\|\theta-\theta(0)\right\|_{2}^{2}\leq R_{\theta}^{2}$. Since $\mathbb{E}_{\text{init}}\left[\left\|\theta_{i}\right\|_{2}^{-2}\right]\leq\infty,\forall i\in[m]$ by the initialization scheme (16), we have that the RHS of (64) is $\mathcal{O}(R_{\theta}m^{1/2})$. Thus, we can obtain
@@
$\mathbb{E}_{\text{init}}\left[\left|f_{\theta}(s,a)-f_{\theta}^{0}(s,a)\right|\right]$ $\leq\mathbb{E}_{\text{init}}\left[G_{1}+G_{2}\right]=\mathcal{O}\left(R_{\vartheta}^{3/2}m^{-1/4}\hat{R}_{\theta}^{1/2}\Delta^{-1/2}+m^{1/2}\Delta^{2}R_{0}^{-1}\right).$
@@
Given model parameter $\vartheta$, which is $\theta$ for the policy and $w$ for the value function, and any pair of actions $a$ and $a^{\prime}$, we have
$\left|u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a^{\prime})\right|$ $=\mathcal{O}\left(R_{\vartheta}\right),$ (66)
$\left|u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a)\right|$ $=\mathcal{O}\left(R_{\vartheta}\right).$ (67)
Proof  By the Jensen’s inequality, we have
@@
$\leq\frac{1}{m}\mathbb{E}_{\text{init}}\left[\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[\vartheta]_{i}^{T}(s,a)>0\right\}[\vartheta]_{i}^{T}(s,a)-\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})>0\right\}[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right|^{2}\right].$
@@
$\leq$ $\frac{1}{m}\mathbb{E}_{\text{init}}\left[\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[\vartheta]_{i}^{T}(s,a)>0\right\}\left([\vartheta]_{i}^{T}(s,a)-[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right)\right.\right.$
$\left.\left.+\sum_{i}^{m}b_{i}\cdot\left(\mathbf{1}\left\{[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})>0\right\}-\mathbf{1}\left\{[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})>0\right\}\right)[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right|^{2}\right]$
$\leq$ $\frac{1}{m}\mathbb{E}_{\text{init}}\left[2\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{[\vartheta]_{i}^{T}(s,a)>0\right\}\left([\vartheta]_{i}^{T}(s,a)-[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right)\right|^{2}\right.$
$\left.+2\left|\sum_{i}^{m}b_{i}\cdot\left(\mathbf{1}\left\{[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})>0\right\}-\mathbf{1}\left\{[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})>0\right\}\right)[\vartheta^{\prime}]_{i}^{T}(s,a^{\prime})\right|^{2}\right].$

### v7

@@ 2.1 Federated Reinforcement Learning
We model the local learning problem of each client as a finite-state infinite-horizon discounted Markov Decision Process (MDP). Accordingly, the FRL system consists of $N$ finite MDPs $\{(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)|n\in\{1,...,N\}\}$, where $\mathcal{S}$ denotes a finite set of states, $\mu$ represents the initial state distribution, $\mathcal{A}$ is a finite set of actions, and $\gamma\in(0,1)$ is the discount factor. The transition function $P_{n}(s^{\prime}|s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ represents the probability that the n-th MDP transits from state $s$ to $s^{\prime}$ after taking action $a$ (Sutton and Barto, 2018). The reward function $\mathcal{R}(s,a):\mathcal{S}\times\mathcal{A}\to[0,R_{\max}]$ gives the expected reward for taking action $a$ in state $s$, and we assume rewards are bounded and non-negative. As a result, the $n$-th MDP $\mathcal{M}_{n}$ can be represented by a 6-tuple $(\mathcal{S},\mu,\mathcal{A},P_{n},\mathcal{R},\gamma)$ sharing the same state space, action space, initial state distribution, and reward function with other clients, but with possibly different transition probabilities. A client’s behavior is controlled by a stochastic policy $\pi_{n}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ which outputs the probability of taking an action $a$ in a given state $s$. Throughout this work, we consider parameterized policies $\pi^{\theta}$ where $\pi^{\theta}(a|s)$ is a differentiable function of the parameter vector $\theta$. For parameterized policies with t-indexed notation, we omit the parameter vector for notation simplicity and write $\pi^{t}$ and $\pi^{t}_{n}$ for the global policy $\pi^{\theta^{t}}$ and the $n$-th local policy $\pi^{\theta^{t}_{n}}$, respectively. Furthermore, we define the state-value function
@@ 3.4 Impact of Aggregation Error
where $m$ is the width of the network, $\theta=(\theta_{i},\dots\theta_{m})^{T}\in\mathcal{R}^{m(d_{s}+d_{a})}$, $w=(w_{i},\dots w_{m})^{T}\in\mathcal{R}^{md_{s}}$ denotes the input weights, and $b_{i}\in\{-1,1\},\forall i\in[m]$ is the output weight. Without loss of generality, we assume that $a\in\mathbb{R}^{d_{a}},s\in\mathbb{R}^{d_{s}}$, $\left\|(s,a)\right\|_{2}\leq 1$, and denote $d=d_{s}+d_{a}$. Given arbitrary $\hat{R}_{\vartheta}<\infty$, we consider the following parameter initialization, where $\vartheta$ is a placeholder for $\theta$ and $w$:
@@
$\mathbb{E}_{\text{init}}\left[\bar{\varepsilon}_{w}\right]$ $=\mathcal{O}\left(R_{w}^{6/5}m^{-1/10}\hat{R}_{w}^{2/5}\right),$ (18)
$\mathbb{E}_{\text{init}}\left[\bar{\varepsilon}_{\theta}\right]$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right),$ (19)
$\mathbb{E}_{\text{init}}\left[\tilde{\varepsilon}_{\theta}\right]$ $=\mathcal{O}\left(R_{\theta}^{1/2}\right).$ (20)
@@ 4.2 Implementation
To compute $\Delta_{n}$, we adopt the tabular maximum likelihood model (Moerland et al., 2020; Strehl et al., 2009; Ornik and Topcu, 2021; Strehl and Littman, 2008) in the implementation of FedPOHCS. Given a set of trajectories, let $C_{n}(s,a)$ denote the number of times action $a$ was taken under state $s$, $C_{n}(s,a,s^{\prime})$ represent the number of times the MDP transited from state $s$ to $s\prime$ after taking action $a$, and $r_{n}$ denote the corresponding sequence of reward received. Generally speaking, the more trajectories we have, the more accurate the modeling will be. For Mountain Cars and Hoppers, the number of trajectories is 200. For HongKongOSMs, the number of trajectories is 1000. We can estimate the transition probability $P_{n}$ of the $n$-th MDP and the reward function $\mathcal{R}$ by
@@ 5 Experiments
In this section, we introduce two federated environments for empirical evaluation and evaluate the proposed client selection method from different perspectives. All experimental results are reproducible and can be accessed from: https://github.com/ShiehShieh/FedPOHCS.
@@ 5.1 Environments
HongKongOSMs consists of 10 equally weighted traffic networks, each based on one of the OSM datasets as shown in Figure 1. Each traffic network contains one RL-controlled and 10 IDM-controlled (Intelligent Driver Model) vehicles. The 18-dimensional observation includes headway, speed, and positional information of visible neighborhoods of the RL-controlled vehicle, and only observations for vehicles running on the same and adjacent lanes are visible to the RL-controlled vehicle. The 2-dimensional action includes acceleration and lane-changing decisions. We adopted the Eclipse Simulation of Urban MObility (SUMO) simulator to conduct this experiment.
Figure 1: OSM datasets of ten areas in Hong Kong. (a) Causeway; (b) Central; (c) Chai Wan; (d) Clear Water Bay; (e) Kennedy Town; (f) Kai Tak; (g) North Point; (h) Po Lam; (i) Sham Shui Po; (j) Tseung Kwan O.
@@ Appendix G Proof of Lemma 12
$u_{w^{t}}^{0}(s)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\left(w^{0}_{i}\right)^{T}(s)>0\right\}\left(w^{t}_{i}\right)^{T}(s),$
$f_{\theta^{t}}^{0}(s,a)=\frac{1}{\sqrt{m}}\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)>0\right\}\left(\theta^{t}_{i}\right)^{T}(s,a).$
@@
For $w\in\mathcal{B}_{R_{w}}^{0},\theta\in\mathcal{B}_{R_{\theta}}^{0},s\in\mathcal{S}$, and $a\in\mathcal{A}$, we have
@@
Proof  Given any pair of model parameters $\theta\in\mathcal{B}_{R_{\theta}}^{0}$ and $\theta^{\prime}\in\mathcal{B}_{R_{\theta}}^{0}$,
$\mathbf{1}\left\{\theta_{i}^{T}(s,a)>0\right\}\neq\mathbf{1}\left\{\left(\theta^{\prime}_{i}\right)^{T}(s,a)>0\right\}$
@@
$\left|\left(\theta^{\prime}_{i}\right)^{T}(s,a)\right|\leq\left|\theta_{i}^{T}(s,a)-\left(\theta^{\prime}_{i}\right)^{T}(s,a)\right|\leq\left\|\theta_{i}-\theta^{\prime}_{i}\right\|_{2}.$
@@
$\left|f_{\theta}(s,a)-f_{\theta}^{0}(s,a)\right|$ $=\frac{1}{\sqrt{m}}\left|\sum_{i=1}^{m}b_{i}\cdot\left(\mathbf{1}\left\{\theta^{T}_{i}(s,a)>0\right\}-\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)>0\right\}\right)\cdot\theta_{i}^{T}(s,a)\right|$
$\leq\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\left|\mathbf{1}\left\{\theta^{T}_{i}(s,a)>0\right\}-\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)>0\right\}\right|\cdot\left|\theta_{i}^{T}(s,a)\right|$
$\leq\frac{1}{\sqrt{m}}\sum_{i=1}^{m}\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}\cdot\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}.$ (60)
@@
  • Case 1: $G_{1}=\frac{1}{\sqrt{m}}\sum_{i\in C_{1}}\left(\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}\right)\cdot\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}$, where $C_{1}=\left\{i\in[m]:\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\leq\Delta\right\}$ and $\Delta>0$.
@@
    $G_{1}$ $\leq\frac{1}{\sqrt{m}}\sum_{i\in C_{1}}\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}=\mathcal{O}\left(m^{1/2}\Delta^{2}R_{0}^{-1}\right).$ (61)
@@
  • Case 2: $G_{2}=\frac{1}{\sqrt{m}}\sum_{i\in C_{2}}\left(\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}\right)\cdot\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}$, where $C_{2}=\left\{i\in[m]:\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}>\Delta\right\}$ and $\Delta>0$.
    We have $\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}/\left\|\theta^{0}_{i}\right\|_{2}>\Delta/\hat{R}_{\theta}$. Consequently, there exits a constant $c\geq\hat{R}_{\theta}/\Delta$ such that for any layer $i\in C_{2},a\in\mathcal{A},s\in\mathcal{S}$, it holds that
    $\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}\leq 1\leq c\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}/\left\|\theta^{0}_{i}\right\|_{2}.$ (62)
    Next, by the Cauchy-Schwarz inequality and $\left\|\theta-\theta^{0}\right\|_{2}\leq R_{\theta}$, we have
    $G_{2}$ $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{\sum_{i\in C_{2}}\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}}.$ (63)
@@
    $\mathbb{E}_{\text{init}}\left[G_{2}\right]$ $\leq\frac{R_{\theta}}{\sqrt{m}}\mathbb{E}_{\text{init}}\left[\sqrt{\sum_{i\in C_{2}}\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}}\right]$
    $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\mathbf{1}\left\{\left(\theta^{0}_{i}\right)^{T}(s,a)\leq\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}\right\}\right]}$
    $\leq\frac{R_{\theta}}{\sqrt{m}}\sqrt{c\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}/\|\theta^{0}_{i}\|_{2}\right]}.$
@@
    $\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}/\|\theta^{0}_{i}\|_{2}\right]$ $\leq\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}^{2}\right]^{1/2}\cdot\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\|\theta^{0}_{i}\|_{2}^{-2}\right]^{1/2}$
    $\leq R_{\theta}\cdot\mathbb{E}_{\text{init}}\left[\sum_{i\in C_{2}}\|\theta^{0}_{i}\|_{2}^{-2}\right]^{1/2},$ (64)
    where the second inequality follows from $\sum_{i=1}^{m}\left\|\theta_{i}-\theta^{0}_{i}\right\|_{2}^{2}=\left\|\theta-\theta^{0}\right\|_{2}^{2}\leq R_{\theta}^{2}$. Since $\mathbb{E}_{\text{init}}\left[\left\|\theta_{i}\right\|_{2}^{-2}\right]\leq\infty,\forall i\in[m]$ by the initialization scheme (16), we have that the RHS of (64) is $\mathcal{O}(R_{\theta}m^{1/2})$. Thus, we can obtain
@@
$\mathbb{E}_{\text{init}}\left[\left|f_{\theta}(s,a)-f_{\theta}^{0}(s,a)\right|\right]$ $\leq\mathbb{E}_{\text{init}}\left[G_{1}+G_{2}\right]=\mathcal{O}\left(R_{\theta}^{3/2}m^{-1/4}\hat{R}_{\theta}^{1/2}\Delta^{-1/2}+m^{1/2}\Delta^{2}R_{0}^{-1}\right).$
@@
For state $s\in\mathcal{S}$, any pair of actions $a$ and $a^{\prime}$, and model parameters $\vartheta,\vartheta^{\prime}\in\mathcal{B}_{R_{\vartheta}}^{0}$, which is $\theta$ for the policy and $w$ for the value function, we have
$\mathbb{E}_{\text{init}}\left[\left|u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a^{\prime})\right|\right]$ $=\mathcal{O}\left(R_{\vartheta}\right),$ (66)
$\mathbb{E}_{\text{init}}\left[\left|u_{\vartheta}(s,a)-u_{\vartheta^{\prime}}(s,a)\right|\right]$ $=\mathcal{O}\left(R_{\vartheta}\right).$ (67)
Proof  By Jensen’s inequality, we have
@@
$\leq\frac{1}{m}\mathbb{E}_{\text{init}}\left[\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\vartheta_{i}^{T}(s,a)>0\right\}\vartheta_{i}^{T}(s,a)-\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})>0\right\}\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right|^{2}\right].$
@@
$\leq$ $\frac{1}{m}\mathbb{E}_{\text{init}}\left[\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\vartheta_{i}^{T}(s,a)>0\right\}\left(\vartheta_{i}^{T}(s,a)-\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right)\right.\right.$
$\left.\left.+\sum_{i}^{m}b_{i}\cdot\left(\mathbf{1}\left\{\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})>0\right\}-\mathbf{1}\left\{\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})>0\right\}\right)\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right|^{2}\right]$
$\leq$ $\frac{1}{m}\mathbb{E}_{\text{init}}\left[2\left|\sum_{i}^{m}b_{i}\cdot\mathbf{1}\left\{\vartheta_{i}^{T}(s,a)>0\right\}\left(\vartheta_{i}^{T}(s,a)-\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right)\right|^{2}\right.$
$\left.+2\left|\sum_{i}^{m}b_{i}\cdot\left(\mathbf{1}\left\{\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})>0\right\}-\mathbf{1}\left\{\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})>0\right\}\right)\left(\vartheta^{\prime}_{i}\right)^{T}(s,a^{\prime})\right|^{2}\right].$